# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
Nonlocal Coulomb interactions play a crucial role in stabilizing distinct electronic phases in kagome materials. In this work, we systematically investigate the effects of on-site ($U$) and inter-site ($V$) Coulomb interactions on the electronic structure and stability of charge-density-wave (CDW) phases in the kagome metal KV3Sb5 using density functional theory (DFT+$U$+$V$) calculations. We demonstrate that $V$ promotes the formation and stability of CDW phases, whereas $U$ suppresses these phases, highlighting a fundamental competition between local and nonlocal electronic correlations. By directly comparing our theoretical results with angle-resolved photoemission spectroscopy (ARPES) data, we identify realistic values of $U$ and $V$ that accurately describe the electronic band structure of KV3Sb5. Our findings establish a detailed $U$$V$ phase diagram for KV3Sb5, offering valuable insights into the correlated electronic states in kagome metals and serving as a foundation for future explorations of correlation-driven phenomena in related materials.
@@ I Introduction
Kagome lattices are formed by a periodic arrangement of corner-sharing triangles, resulting in a distinctive crystal structure with inherent geometric frustration Freedman et al. (2010). This geometric frustration arises from the nonequivalence of the three vertices in each triangular unit, which prevents the system from achieving long-range magnetic ordering and instead stabilizes disordered spin states. Consequently, kagome lattices with intrinsic magnetic frustration are promising candidates for hosting quantum spin liquid states Balents (2010); Norman (2016); Georgopoulou et al. (2023); Han et al. (2012).
In addition, kagome lattices exhibit topologically nontrivial electronic states linked to local electronic symmetry breaking Ye et al. (2018); Yin et al. (2018); Liu et al. (2020); Kang et al. (2020); Morali et al. (2019). Their complex band structures—featuring flat bands, Dirac cones, and van Hove singularities—heighten their significance in condensed matter physics Yin et al. (2022). The interplay between strong Coulomb interactions and nontrivial band topology, driven by the diverging density of states at van Hove singularities and flat bands, gives rise to diverse exotic phenomena. These include topological insulator Guo and Franz (2009), superconductivity Ko et al. (2009), charge density wave (CDW) Wang et al. (2013), charge fractionalization Rüegg and Fiete (2011); Mizoguchi et al. (2017), and Dirac/Weyl semimetals Morali et al. (2019); Liu et al. (2019).
@@
Regarding their crystal structures, initial results from X-ray diffraction Ortiz et al. (2020, 2021b) and scanning tunneling microscopy (STM) Jiang et al. (2021); Zhao et al. (2021); Liang et al. (2021a) indicate the presence of a $2\times 2$ CDW phase. A recent density functional theory (DFT) study, in comparison with STM data, confirmed the inverse Star of David (ISD) distortion as the ground state for AV3Sb5 compounds Tan et al. (2021). Subsequent studies have provided further evidence supporting the ISD deformation in these systems Kato et al. (2022); Kang et al. (2022).
Important features of the electronic structures in $A$V3Sb5 compounds are the van Hove singularities near the Fermi energy, which have V-$3d$ orbital character. The precise position of these van Hove singularities depends on details of the Coulomb interactions, inducing electronic instabilities that lead to the emergence of various CDW phases Jiang et al. (2021); Denner et al. (2021); Jeong et al. (2022). In particular, the chiral CDW phase observed in $A$V3Sb5 has been plausibly attributed to nonlocal Coulomb interactions Denner et al. (2021). Other contributing factors, such as lattice deformations, electron-phonon coupling, and Fermi surface nesting, can also interplay with electronic correlations and play significant roles in mediating the CDW phase transition Tan et al. (2021); Wenzel et al. (2022); Xie et al. (2022); Ratcliff et al. (2021); Luo et al. (2022); Wu et al. (2022); Liu et al. (2022).
The impact of Coulomb interactions has been extensively investigated using Kagome Hubbard model calculations, primarily within single-band models. Kiesel et al. Kiesel and Thomale (2012); Kiesel et al. (2013) explored the system using a single-band extended Hubbard model that incorporates both on-site ($U$) and inter-site ($V$) Coulomb interactions. They studied various phases arising from Fermi surface instabilities within the $U$-$V$ phase diagram Kiesel and Thomale (2012); Kiesel et al. (2013). Notably, near the van Hove filling, $V$ can promote distinct CDW phases, which explain several key experimental findings Ferrari et al. (2022). The presence of sizable long-range interactions has been also reported, which exhibit interesting universality across kagome metals Di Sante et al. (2023). More recent studies have further emphasized the critical role of inter-site Coulomb interactions in understanding the electronic structure of kagome materials, investigating their impact on a broad range of electronic properties Wu et al. (2021); Zheng et al. (2022); Kiran et al. (2024); Jiang et al. (2025); Dong et al. (2023); Bigi et al. (2024); Wenger et al. (2024).
@@ II Computational details
Density functional theory (DFT) calculations were performed using the plane-wave pseudopotential method, as implemented in the Quantum Espresso package Giannozzi et al. (2009, 2017a, 2017b). The exchange-correlation functional was treated using the generalized gradient approximation (GGA) parameterized by Perdew-Burke-Ernzerhof (PBE) Perdew et al. (1996). Pseudopotentials were obtained from the Standard Solid-State Pseudopotentials (SSSP) library, version 1.2.1 SSS . We considered a $2\times 2\times 1$ supercell of KV3Sb5, containing 36 atoms, to incorporate two relevant CDW phases--SD and ISD--on equal footing (see Fig. 1), and performed all calculations using a $5\times 5\times 5$ $k$-mesh. The plane-wave energy cut-off was set to 37 Ry. For density of states calculations, a denser $8\times 8\times 8$ $k$-mesh was used. Band unfolding calculations were carried out using the unfold.x code Pacilè et al. (2021), based on the method of Popescu and Zunger Popescu and Zunger (2012). Spin-orbit coupling and van der Waals corrections were not included, as they do not affect the main conclusions of this study.
The optimized structure obtained from the DFT calculations yields lattice parameters of $a=5.48$ and $c=9.35$. These optimized lattice parameters are in good agreement with experimental data Ortiz et al. (2019, 2021b) and previous DFT studies Tan et al. (2021); Ortiz et al. (2021b). Neglecting van der Waals corrections results in a slightly larger $c$ lattice parameter compared to experimental reports ($a_{\text{expt}}=5.48$, $c_{\text{expt}}=8.958$) Ortiz et al. (2021b, 2019). Details of the structural parameters are provided in Table S1 of the Supplementary Material Sup .
The Coulomb interaction parameters ($U$ and $V$) for the V-$3d$ manifold were calculated using the constrained random-phase approximation (cRPA) with the projector method Kaltak (2015) implemented in VASP Kresse and Furthmüller (1996). A $8\times 8\times 5$ $k$-mesh and 288 bands were employed for the cRPA calculations. To construct the Wannier functions, all V-$3d$ and Sb-$5p$ orbitals were considered, with a disentanglement energy window of [-5.0, 4.8] eV (see Fig. S1). The Wannier functions were obtained using WANNIER90 Mostofi et al. (2008) and the VASP2WANNIER Franchini et al. (2012) interface, with default parameters Reddy et al. (2024) for disentanglement and Wannierization. A large plane-wave cutoff energy of 500 eV was used for self-consistent calculations, while 400 eV was used for dielectric function calculations. The resulting band structures and interaction parameters are provided in the Supplementary Material Sup . Using the $U$ and $V$ values obtained from cRPA calculations, we employed the DFT+$U$+$V$ method, as implemented in Quantum Espresso Campo and Cococcioni (2010), to investigate the electronic structure of KV3Sb5. In addition, we systematically varied the values of $U$ and $V$ to further investigate their impact on the ground state and electronic structure of KV3Sb5.
@@ III.1 Pristine and CDW structures
Figures 1(a) and (b) show the unit cell and $2\times 2$ supercell of KV3Sb5, respectively, with space group P6/$mmm$ (No. 191). The supercell contains two inequivalent vanadium sites, denoted as V1 and V2, whereas the unit cell contains only one. The vanadium sublattice forms a perfect kagome lattice structure, with the V-Sb layers inserted between the K atoms. Two different $2\times 2$ CDW structures, SD (Fig. 1(c)) and ISD (Fig. 1(d)), were constructed by incorporating opposite lattice distortions on the V atoms relative to the pristine supercell.
III.2 Stability of CDW phases depending on $U$ and $V$
@@
First, we discuss the stability of the CDW phases upon introducing $U$ and $V$ in the kagome metal KV3Sb5. Figure 2 shows the energy profiles as a function of vanadium atomic distortions for different values of $U$ and $V$. Positive (negative) distortion values correspond to the ISD (SD) phase. Our results clearly demonstrate that, in the absence of $U$ and $V$, the ISD deformation is energetically more favorable than the SD deformation. This is consistent with experimental STM images and previous DFT results, where the ISD phase is found to be more stable than the SD phase Tan et al. (2021); Jiang et al. (2021). For all fixed values of $U$, increasing $V$ not only enhances the stability of both the ISD and SD phases but also increases atomic distortions (see Figs. 2(a-d)). The ISD phase remains more stable than the SD phase upon variation of $V$. For instance, at $U$ = 0, the atomic distortion increases by a factor of 2.9 (ISD phase) and 3.2 (SD phase) for $V=3$ eV compared to the case of $V=0$. Furthermore, the energy difference between the ISD and SD phases reaches 204 meV, a significant increase from 8 meV at $U$ = 0 and $V=0$.
Interestingly, for a fixed $V$, increasing $U$ progressively reduces the stability of both the ISD and SD phases. Eventually, at $V=0$ and $U=3$ eV, the pristine phase becomes energetically more favorable than the CDW phases (see Fig. 2(e)). This demonstrates that strong local Coulomb interaction alone cannot stabilize either of the CDW phases. For finite values of $V$, higher $U$ values ($\geq 3$ eV) are required to stabilize the pristine phase relative to the CDW phases. These results indicate that $V$ and $U$ exert opposite effects: while $V$ stabilizes CDW phases, $U$ suppresses them. This contrast provides compelling evidence that incorporating both $U$ and $V$ leads to fundamentally different physical consequences in the kagome metal KV3Sb5.
We compare these findings for KV3Sb5 with those for another kagome metal, CsV3Sb5. Employing the similar approaches, we obtained the corresponding energy profile for CsV3Sb5 as shown in Fig. S2 of the Supplementary Material Sup . For a fixed $U$, CsV3Sb5 follows a similar trend to KV3Sb5, except when $V=0$. When $V=0$, as $U$ increases (for $U<3$ eV), the ISD phase gradually destabilizes prior to the SD phase. At $U=3$ eV, the SD phase also becomes destabilized, approaching the pristine phase. These results shows common tendency toward ISD phase for both systems. Note that the effective values of $U$ and $V$ in CsV3Sb5 are smaller than in KV3Sb5 suggesting the stronger CDW instability of the system.
@@
III.3 Phase diagram of KV3Sb5
Previous studies on the $A$V3Sb5 system have demonstrated the possibility of various CDW phases, which not only differentiate in-plane SD and ISD configurations but also include phase shifts between adjacent kagome layers Liang et al. (2021a); Li et al. (2021); Ortiz et al. (2021a); Tan et al. (2021) (see Fig. S3). Motivated by these findings, we considered six distinct CDW phases by employing the SD and ISD deformations with either zero or $\pi$ phase shifts between adjacent kagome layers in a $2\times 2\times 2$ supercell. These six CDW phases include (a) ISD+ISD ($\pi$-shift), (b) ISD+ISD (zero-shift), (c) SD+SD ($\pi$-shift), (d) SD+SD (zero-shift), (e) SD+ISD (zero-shift), and (f) SD+ISD ($\pi$-shift). Note that the ISD+ISD with zero phase shift is equivalent to the $2\times 2\times 1$ ISD structure discussed in Sec. III.2.
Figure 3 shows the overall CDW phase diagram of KV3Sb5 as a function of $U$ and $V$. We first note that within the studied $U$-$V$ parameter range, the ISD deformation is consistently favored, as shown in Fig. 2. This preference arises because the energy scale of the inter-layer coupling is significantly smaller than that of the intra-layer CDW distortion. Except in the limit of large $U$, where the pristine phase is stabilized for small $V$, ISD-related phases remain more stable than SD phases throughout the phase diagram.
Among the various phases, two dominant ISD-type CDW phases, ISD+ISD ($\pi$-shift) and ISD+ISD (zero-shift), are noted in the phase diagram. We found that in the small $U$ and $V$ regime, ISD+ISD ($\pi$-shift) is stabilized, as shown in Fig. 3. We find that inter-site Coulomb interactions promote coherent ISD deformations across different kagome layers, favoring a zero-shift phase alignment. For sufficiently large $V$, ISD+ISD (zero-shift) is always stabilized regardless of $U$. However, in the small $V$ regime, competition arises between ISD+ISD ($\pi$-shift) and either ISD+ISD (zero-shift) or the pristine phase, depending on whether $U$ is small or large. This finding is intriguing because our Coulomb interaction profiles are restricted to the in-plane interactions, yet they selectively stabilize specific inter-layer phase shifts. It also highlights that competing phases can be stabilized by subtle details of the electronic structure.
Previous studies have identified the ISD+ISD ($\pi$-shift) phase as the ground state of the $A$V3Sb5 system Liang et al. (2021b); Jin et al. (2024); Li et al. (2022). Hence, we can determine the realistic regime of the on-site $U$ and inter-site $V$ Hubbard parameters for the KV3Sb5 system in the $U$-$V$ phase diagram: small to moderate $U$ values and small $V$ values (see the blue-filled areas in Fig. 3). To further quantify the Coumonb interaction parameters, we performed cRPA calculations, obtaining $U$ = 1.35 eV and $V$ = 0.17 eV. The corresponding point is marked with a star symbol in Fig. 3, demonstrating that KV3Sb5 lies well within the ISD+ISD ($\pi$-shift) regime, consistent with experimental findings. The validity of these $U$ and $V$ values obtained from cRPA will be further examined in the next section through direct comparison between the band dispersion from DFT+$U$+$V$ calculations and experimental ARPES data.
III.4 Electronic structures: band structure

### v2

@@ Abstract
Nonlocal Coulomb interactions play a crucial role in stabilizing distinct electronic phases in kagome materials. In this work, we systematically investigate the effects of on-site ($U$) and inter-site ($V$) Coulomb interactions on the electronic structure and stability of charge-density-wave (CDW) phases in the kagome metal KV3Sb5 using density functional theory (DFT+UU+$V$) calculations. We demonstrate that $V$ promotes the formation and stability of CDW phases, whereas $U$ suppresses these phases, highlighting a fundamental competition between local and nonlocal Coulomb interactions. By directly comparing our theoretical results with angle-resolved photoemission spectroscopy (ARPES) data, we identify realistic values of $U$ and $V$ that accurately describe the electronic band structure of KV3Sb5. Our findings establish a detailed UUVV phase diagram for KV3Sb5, offering valuable insights into the correlated electronic states in kagome metals and serving as a foundation for future explorations of correlation-driven phenomena in related materials.
@@ I Introduction
Kagome lattices are formed by a periodic arrangement of corner-sharing triangles, resulting in a distinctive crystal structure with inherent geometric frustration [1]. This geometric frustration arises from the nonequivalence of the three vertices in each triangular unit, which prevents the system from achieving long-range magnetic ordering and instead stabilizes disordered spin states. Consequently, kagome lattices with intrinsic magnetic frustration are promising candidates for hosting quantum spin liquid states [2, 3, 4, 5].
In addition, kagome lattices exhibit topologically nontrivial electronic states linked to local electronic symmetry breaking [6, 7, 8, 9, 10]. Their complex band structures—featuring flat bands, Dirac cones, and van Hove singularities—heighten their significance in condensed matter physics [11]. The interplay between strong Coulomb interactions and nontrivial band topology, driven by the diverging density of states at van Hove singularities and flat bands, gives rise to diverse exotic phenomena. These include topological insulator [12], superconductivity [13], charge density wave (CDW) [14], charge fractionalization [15, 16], and Dirac/Weyl semimetals [10, 17].
@@
Regarding their crystal structures, initial results from X-ray diffraction [27, 28] and scanning tunneling microscopy (STM) [21, 22, 23] indicate the presence of a $2\times 2$ CDW phase. A recent density functional theory (DFT) study, in comparison with STM data, confirmed the inverse Star of David (ISD) distortion as the ground state for AV3Sb5 compounds [36]. Subsequent studies have provided further evidence supporting the ISD deformation in these systems [37, 38].
Important features of the electronic structures in AAV3Sb5 compounds are the van Hove singularities near the Fermi energy, which have V-$3d$ orbital character. The precise position of these van Hove singularities depends on details of the Coulomb interactions, inducing electronic instabilities that lead to the emergence of various CDW phases [21, 39, 40]. In particular, the chiral CDW phase observed in AAV3Sb5 has been plausibly attributed to nonlocal Coulomb interactions [39]. Other contributing factors, such as lattice deformations, electron-phonon coupling, and Fermi surface nesting, can also interplay with electronic correlations and play significant roles in mediating the CDW phase transition [36, 41, 42, 43, 44, 45, 46].
The impact of Coulomb interactions has been extensively investigated using Kagome Hubbard model calculations, primarily within single-band models. Kiesel et al. [47, 48] explored the system using a single-band extended Hubbard model that incorporates both on-site ($U$) and inter-site ($V$) Coulomb interactions. They studied various phases arising from Fermi surface instabilities within the $U$-$V$ phase diagram [47, 48]. Notably, near the van Hove filling, $V$ can promote distinct CDW phases, which explain several key experimental findings [49]. The presence of sizable long-range interactions has been also reported, which exhibit interesting universality across kagome metals [50]. More recent studies have further emphasized the critical role of inter-site Coulomb interactions in understanding the electronic structure of kagome materials, investigating their impact on a broad range of electronic properties [51, 52, 53, 54, 55, 56, 57].
@@ II Computational details
Density functional theory (DFT) calculations were performed using the plane-wave pseudopotential method, as implemented in the Quantum Espresso package [58, 59, 60]. The exchange-correlation functional was treated using the generalized gradient approximation (GGA) parameterized by Perdew-Burke-Ernzerhof (PBE) [61]. Pseudopotentials were obtained from the Standard Solid-State Pseudopotentials (SSSP) library, version 1.2.1 [62]. We considered a $2\times 2\times 1$ supercell of KV3Sb5, containing 36 atoms, to incorporate two relevant CDW phases-SD and ISD-on equal footing (see Fig. 1), and performed all calculations using a $5\times 5\times 5$ $k$-mesh. The plane-wave energy cut-off was set to 37 Ry. For density of states calculations, a denser $8\times 8\times 8$ $k$-mesh was used. Band unfolding calculations were carried out using the unfold.x code [63], based on the method of Popescu and Zunger [64]. Spin-orbit coupling and van der Waals corrections were not included, as they do not affect the main conclusions of this study.
The optimized structure obtained from the DFT calculations yields lattice parameters of $a=5.48$ Å and $c=9.35$ Å. These optimized lattice parameters are in good agreement with experimental data [18, 28] and previous DFT studies [36, 28]. Neglecting van der Waals corrections results in a slightly larger $c$ lattice parameter compared to experimental reports ($a_{\text{expt}}=5.48$ Å, $c_{\text{expt}}=8.958$ Å) [28, 18]. Details of the structural parameters are provided in Table S1 of the Supplementary Material [65].
The Coulomb interaction parameters ($U$ and $V$) for the V-$3d$ manifold were calculated using the constrained random-phase approximation (cRPA) with the projector method [66] implemented in VASP [67]. A $8\times 8\times 5$ $k$-mesh and 288 bands were employed for the cRPA calculations. To construct the Wannier functions, all V-$3d$ and Sb-$5p$ orbitals were considered, with a disentanglement energy window of [-5.0, 4.8] eV (see Fig. S1). The Wannier functions were obtained using WANNIER90 [68] and the VASP2WANNIER [69] interface, with default parameters [70] for disentanglement and Wannierization. A large plane-wave cutoff energy of 500 eV was used for self-consistent calculations, while 400 eV was used for dielectric function calculations. The resulting band structures and interaction parameters are provided in the Supplementary Material [65]. Using the $U$ and $V$ values obtained from cRPA calculations, we employed the DFT+UU+$V$ method, as implemented in Quantum Espresso [71], to investigate the electronic structure of KV3Sb5. In addition, we systematically varied the values of $U$ and $V$ to further investigate their impact on the ground state and electronic structure of KV3Sb5.
@@ III.2 Pristine and CDW structures
Figures 1(a) and (b) show the unit cell and $2\times 2$ supercell of KV3Sb5, respectively, with space group P6/mmmmmm (No. 191). The supercell contains two inequivalent vanadium sites, denoted as V1 and V2, whereas the unit cell contains only one. The distinct Wyckoff positions for V1 and V2 are provided in Table S1 of the Supplementary Material [65]. The vanadium sublattice forms a perfect kagome lattice structure, with the V-Sb layers inserted between the K atoms. Two different $2\times 2$ CDW structures, SD (Fig. 1(c)) and ISD (Fig. 1(d)), were constructed by incorporating opposite lattice distortions on the V atoms relative to the pristine supercell.
III.3 Stability of CDW phases depending on $U$ and $V$
@@
First, we discuss the stability of the CDW phases upon introducing $U$ and $V$ in the kagome metal KV3Sb5. Figure 2 shows the energy profiles as a function of vanadium atomic distortions for different values of $U$ and $V$. Positive (negative) distortion values correspond to the ISD (SD) phase. Our results clearly demonstrate that, in the absence of $U$ and $V$, the ISD deformation is energetically more favorable than the SD deformation. This is consistent with experimental STM images and previous DFT results, where the ISD phase is found to be more stable than the SD phase [36, 21]. For all fixed values of $U$, increasing $V$ not only enhances the stability of both the ISD and SD phases but also increases atomic distortions (see Figs. 2(a-d)). The ISD phase remains more stable than the SD phase upon variation of $V$. For instance, at $U$ = 0, the atomic distortion increases by a factor of 2.9 (ISD phase) and 3.2 (SD phase) for $V=3$ eV compared to the case of $V=0$. Furthermore, the energy difference between the ISD and SD phases reaches 204 meV, a significant increase from 8 meV at $U$ = 0 and $V=0$. Interestingly, for a fixed $V$, increasing $U$ progressively reduces the stability of both the ISD and SD phases. Eventually, at $V=0$ and $U=3$ eV, the pristine phase becomes energetically more favorable than the CDW phases (see Fig. 2(e)). This demonstrates that strong local Coulomb interaction alone cannot stabilize either of the CDW phases. For finite values of $V$, higher $U$ values ($\geq 3$ eV) are required to stabilize the pristine phase relative to the CDW phases. These results indicate that $V$ and $U$ exert opposite effects: while $V$ stabilizes CDW phases, $U$ suppresses them. This contrast provides compelling evidence that incorporating both $U$ and $V$ leads to fundamentally different physical consequences in the kagome metal KV3Sb5. We compare these findings for KV3Sb5 with those for another kagome metal, CsV3Sb5. Employing the similar approaches, we obtained the corresponding energy profile for CsV3Sb5 as shown in Fig. S2 of the Supplementary Material [65]. For a fixed $U$, CsV3Sb5 follows a similar trend to KV3Sb5, except when $V=0$. When $V=0$, as $U$ increases (for $U<3$ eV), the ISD phase gradually destabilizes prior to the SD phase. At $U=3$ eV, the SD phase also becomes destabilized, approaching the pristine phase. These results shows common tendency toward ISD phase for both systems. Note that the effective values of $U$ and $V$ in CsV3Sb5 are smaller than in KV3Sb5 suggesting the stronger CDW instability of the system.
@@
III.4 Phase diagram of KV3Sb5
Previous studies on the AAV3Sb5 system have demonstrated the possibility of various CDW phases, which not only differentiate in-plane SD and ISD configurations but also include phase shifts between adjacent kagome layers [23, 25, 26, 36] (see Fig. S3). Motivated by these findings, we considered six distinct CDW phases by employing the SD and ISD deformations with either zero or $\pi$ phase shifts between adjacent kagome layers in a $2\times 2\times 2$ supercell. These six CDW phases include (a) ISD+ISD ($\pi$-shift), (b) ISD+ISD (zero-shift), (c) SD+SD ($\pi$-shift), (d) SD+SD (zero-shift), (e) SD+ISD (zero-shift), and (f) SD+ISD ($\pi$-shift). Note that the ISD+ISD with zero phase shift is equivalent to the $2\times 2\times 1$ ISD structure discussed in Sec. III.3. Figure 3 shows the overall CDW phase diagram of KV3Sb5 as a function of $U$ and $V$. We first note that within the studied $U$-$V$ parameter range, the ISD deformation is consistently favored, as shown in Fig. 2. This preference arises because the energy scale of the inter-layer coupling is significantly smaller than that of the intra-layer CDW distortion. Except in the limit of large $U$, where the pristine phase is stabilized for small $V$, ISD-related phases remain more stable than SD phases throughout the phase diagram. Among the various phases, two dominant ISD-type CDW phases, ISD+ISD ($\pi$-shift) and ISD+ISD (zero-shift), are noted in the phase diagram. We found that in the small $U$ and $V$ regime, ISD+ISD ($\pi$-shift) is stabilized, as shown in Fig. 3. We find that inter-site Coulomb interactions promote coherent ISD deformations across different kagome layers, favoring a zero-shift phase alignment. For sufficiently large $V$, ISD+ISD (zero-shift) is always stabilized regardless of $U$. However, in the small $V$ regime, competition arises between ISD+ISD ($\pi$-shift) and either ISD+ISD (zero-shift) or the pristine phase, depending on whether $U$ is small or large. This finding is intriguing because our Coulomb interaction profiles are restricted to the in-plane interactions, yet they selectively stabilize specific inter-layer phase shifts. It also highlights that competing phases can be stabilized by subtle details of the electronic structure. Previous studies have identified the ISD+ISD ($\pi$-shift) phase as the ground state of the AAV3Sb5 system [23, 72, 73]. Hence, we can determine the realistic regime of the on-site $U$ and inter-site $V$ Hubbard parameters for the KV3Sb5 system in the $U$-$V$ phase diagram: small to moderate $U$ values and small $V$ values (see the blue-filled areas in Fig. 3). To further quantify the Coumonb interaction parameters, we performed cRPA calculations, obtaining $U$ = 1.35 eV and $V$ = 0.17 eV. The corresponding point is marked with a star symbol in Fig. 3, demonstrating that KV3Sb5 lies well within the ISD+ISD ($\pi$-shift) regime, consistent with experimental findings. The validity of these $U$ and $V$ values obtained from cRPA will be further examined in the next section through direct comparison between the band dispersion from DFT+UU+$V$ calculations and experimental ARPES data.
III.5 Electronic structures: band structure